OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = 5x + 10; (2, 14)

OpenStudy (anonymous):

use the equation y-y1=m(x-x1) y-14=5(x-2) y-14=5x-10 y=5x+4 The answer is: y=5x+4

OpenStudy (lgbasallote):

the slope of your equation is 5...since you're looking for a parallell line that line is also 5..now use the formula \[\large y -y_1 = m(x- x_1)\]